OpenStudy (anonymous):

Adult tickets for a play cost 25$ each and children's cost 15$ each. The total receits were 7200$ and the total attendence was 400. How many adults and how many children attended?

OpenStudy (anonymous):

a + c = 400 25a+ 15c = 7200

OpenStudy (anonymous):

solve system by substitution or elimination...are you familiar with how to do that?

OpenStudy (anonymous):

a=400-c 25(400-c)+15c = 7200 10000-25c+15c=7200 10000-10c=7200 -10c=-2800 c=280 a=400-280=120

OpenStudy (anonymous):

check 120+280=400 25(120)+15(280)=7200 3000+4200=7200 7200=7200